Select three fans randomly at a football game in which Penn State is playing over Notre Dame. Identify whether the fan is Penn State fan (P) or a Notre Dame fan (D). Let X be the number of Penn State fans.


1
Expert's answer
2021-01-20T13:10:21-0500

The fan is a Penn State fan "(P)" or a Notre Dame fan "(N)." This experiment yields the following sample space:

"S=\{PPP, PPN,PNP, PNN, NPP, NPN, NNP, NNN\}"

Let "X=" the number of Penn State fans selected. The possible values of "X" are, therefore, either "0,1,2," or "3."

Since the game is a home game, let's suppose that 80% of the fans attending the game are Penn State fans, while 20% are Notre Dame fans. That is, "P(P)=0.8" and "P(N)=0.2"

Then, by independence:


"P(X=0)=P(NNN)=(0.2)^3=0.008"

By independence and mutual exclusivity of "NNP, NPN," and "PNN"


"P(X=1)=P(NNP)+P(NPN)+P(PNN)"

"=3(0.8)(0.2)^2=0.096"

By independence and mutual exclusivity of "PPN, PNP," and "NPP"


"P(X=2)=P(PPN)+P(PNP)+P(NPP)"

"=3(0.8)^2(0.2)=0.384"

By independence:


"P(X=3)=P(PPP)=(0.8)^3=0.512"


Check


"0.008+0.096+0.384+0.512=1"

The probability of the sample space is "1."

Because the values that it takes on are random, the variable "X" is a random variable (rv).